OpenStudy (anonymous):

What is the value of 3[29-(9+6)]-4•7 ? MEDEL AND I WILL FAN!!!!!

OpenStudy (emogirlatemoooocow):

It's an improperly-formed expression, missing a closing parenthesis. If you mean 3(29-(9+6))-4*7 then by PEMDAS: 3(29-(9+6))-4*7 = 3(29-15)-4*7 ... evaluate inner parenthesis contents = 3(14)-4*7 ... evaluate outer parenthesis contents = 42-28 ... do multiplications = 14 ... do subtraction

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Help This is the problem

OpenStudy (emogirlatemoooocow):

Want we to walk you through this problem ?

OpenStudy (mrcoolguy):

-79.7

OpenStudy (emogirlatemoooocow):

its using PEMDAS

OpenStudy (emogirlatemoooocow):

so start with [29-(9+6)]